# Unlocking Personalized Knowledge in Federated Large Language Model: The Power of Mixture of Experts

### v1

@@ 5 Experiments
In this section, we will use substantial experiments to demonstrate the effectiveness of FLEx.
@@ 5.1 Experimental Setup
To simulate real-world scenarios, we utilize the Databricks-dolly-15k [^1] https://huggingface.co/datasets/databricks/databricks-dolly-15k dataset, an open-source collection of instruction follow-up records generated by numerous Databricks employees. This dataset covers a variety of behavioral categories as described in the InstructGPT paper ouyangTrainingLanguageModels2022 and includes eight subtasks. We select four subtasks—classification, closed QA, information extraction, and summarization—to create a non-IID scenario. These tasks have fixed responses, unlike open QA or creative writing tasks. Therefore, ROUGE-L serves as an effective evaluation metric to assess the instruction followability capability of large language models.
Implementation.
We utilize the OpenFedLLM framework yeOpenFedLLMTrainingLarge2024 to evaluate federated learning methods, including FedAvg mcmahanCommunicationefficientLearningDeep2017 , FedProx liFederatedOptimizationHeterogeneous2020 , SCAFFOLD karimireddyScaffoldStochasticControlled2020 , FedAvgM hsuMeasuringEffectsNonIdentical2019 , FedAdagrad reddiAdaptiveFederatedOptimization2021 , FedYogi reddiAdaptiveFederatedOptimization2021 and FedAdam reddiAdaptiveFederatedOptimization2021 . Each client trains on 20 samples per round for 100 rounds. We select the same round checkpoint for each client to compare ROUGE-L scores. The Qwen1.5-MoE-A2.7B model [^2] https://huggingface.co/Qwen/Qwen1.5-MoE-A2.7B, configured with bfloat16, serves as the base model and is trained on NVIDIA H20 GPUs with 96GB of memory. We apply LoRA to the attention layer with a rank of 32 and an alpha of 64. The Adam optimizer is used to train the LoRA adapter with a batch size of 4, using a consistent learning rate of 4e-5. The maximum input sequence length is set to 2048, and we provide the Alpaca template as a prompt.
Table 1: Rouge-L performance comparison of different federated learning strategies, including our proposed method, on the Databricks-dolly-15k dataset under a non-IID scenario (one subtask per client). The results are reported for four distinct instruction-based tasks: classification, closed QA, information extraction, and summarization, using the Rouge-L metric to evaluate instruction followability. The best result over federated method is bolded and the second-best result is underlined. CLF: Classification, CQA: Closed Question Answering, IE: Information Extraction, Summ: Summarization
@@
FLEx (Ours) 52.10 38.23 40.13 42.09 43.13 49.74
5.2 Performance under Pathological Non-IID Distribution
We first consider pathological data heterogeneity, where we assign one subtask per client, with a total of four clients. As shown in Table 1, we compared our methods with baseline methods. From this table, we can observe that FLEx (Ours) has the highest average performance (43.13), outperforming all other methods. From an averaging perspective, MoE+FedAvgM (42.37) is the next closest competitor. Also, performance seems to vary by task, except summarization task, FLEx achieves the highest score. From this table, we can observe that the proposed method consistently leads in average performance, indicating its effectiveness across various tasks. Besides, it can be observed that due to the powerful pretrained model, simply apply averaging as traditional federated paradigm will easily cannot outperform simple local update and only our method and MoE+FedAvgM can ourperform local update.
We also evaluated the performance of each client’s checkpoint on the MMLU-STEM benchmark wangMMLUPRO to assess the extent of knowledge forgetting after personalization. Specifically, each client’s model was tested individually, and the average performance was calculated. The results indicate that FLEx (Ours) preserves the most general knowledge while maintaining strong performance in personalization. In contrast, MoE+FedAvgM, which showed competitive results in terms of personalization, demonstrated the greatest loss of general knowledge, highlighting its tendency to forget important shared knowledge during the personalization process. This further demonstrates the effectiveness of our FLEx method, which strikes a better balance between general knowledge retention and personalization.
Table 2: Rouge-L performance comparison of different federated learning strategies, including our proposed method, on the Alpaca-gpt4, Finance-Alpaca and MedAlpaca dataset under a non-IID scenario (one subtask per client). The results are using the Rouge-L metric to evaluate instruction followability. The best result over federated method is bolded and the second-best result is underlined.
Method Alpaca-gpt4 Finance-Alpaca MedAlpaca Avg
Local Training 31.31 28.64 30.93 30.29
MoE + FedAvg 30.11 28.17 30.06 29.44
MoE + FedAvgM 28.90 27.18 28.78 28.28
MoE + FedAdam 29.23 28.70 31.18 27.70
MoE + FedAdagrad 30.24 28.79 30.60 29.87
MoE + SCAFFOLD 29.96 28.38 28.92 29.08
MoE + FedProx 30.05 28.60 30.01 29.55
MoE + FedYogi 29.73 28.42 30.51 29.55
FLEx (Ours) 31.54 29.89 31.31 30.91
Table 2 presents the ROUGE-L performance comparison of different federated learning strategies, including our proposed method, FLEx, on datasets that simulate real-world user scenarios: Alpaca-gpt4, Finance-Alpaca, and MedAlpaca. The experiment was conducted in a highly non-IID setting, where each client was assigned a domain-specific dataset.
FLEx outperforms all other methods across all datasets, achieving the highest ROUGE-L scores: 31.54 for Alpaca-gpt4, 29.89 for Finance-Alpaca, and 31.31 for MedAlpaca, with an average score of 30.91. This demonstrates its robustness in handling diverse and specialized data. The second-best performing method, MoE + FedAdagrad, shows an average score of 29.87, but its performance is inconsistent. While it ranks second on MedAlpaca (31.18), its performance on Alpaca-gpt4 and Finance-Alpaca is lower, highlighting its instability. Interestingly, federated averaging methods (e.g., MoE + FedAvg) and other federated optimizers (MoE + FedAvgM, MoE + FedAdam) fail to outperform the local update method (Local), which achieves an average score of 30.29. This emphasizes that traditional federated learning approaches struggle to effectively aggregate models in highly non-IID scenarios, sometimes even underperforming compared to isolated local training.
These results reinforce the rationale behind our choice of domain-specific datasets like Finance-Alpaca and MedAlpaca, which simulate the specialized needs of real-world users. FLEx excels in such complex, heterogeneous data environments by capturing personalized knowledge and effectively leveraging domain-specific features, outperforming both baseline methods and local training.
5.3 Performance under Dirichlet Non-IID Distribution
Compared to previous experiments on pathological non-IID scenerios, we introduce dirichlet non-IID distributions into our experiments as shown in Table 3. In federated learning (FL), the Dirichlet non-IID distribution refers to the scenario where data is non-identically and independently distributed across clients, with each client’s data distribution being generated from a Dirichlet distribution, typically used to model the unevenness of data distribution in real-world scenarios. Here, we choose $\alpha=1.0$ as our dirichlet distribution parameter.
We introduced the Dirichlet non-IID distribution to further evaluate the robustness of our proposed FLEx method in a more realistic and challenging scenario. The Dirichlet distribution, with its ability to model the unevenness of data distribution across clients, is commonly used in federated learning to simulate real-world data heterogeneity. By increasing the number of clients to 10 in this experiment, we aimed to simulate a more complex and distributed environment, which is closer to practical deployment scenarios in federated learning.
Table 3: ROUGE-L performance comparison of different federated learning strategies, including our proposed method, on the Databricks-dolly-15k dataset under a Dirichlet distribution with 10 clients, aggregating 4 clients per communication round. The results are reported for four distinct instruction-based tasks: classification, closed QA, information extraction, and summarization, using the ROUGE-L metric to evaluate instruction following ability.
@@ 5.3 Performance under Dirichlet Non-IID Distribution
FLEx (Ours) 37.30 37.54
To ensure fairness and consistency, we selected a fixed communication round and evaluated the performance using the aggregated checkpoints of four clients per round, chosen based on their data distributions. For each task, the client whose data distribution most closely matches the task’s focus was selected. For example, if a client’s dataset contained the majority of samples for a specific task, the evaluation for that task was conducted based on the corresponding client’s model. This approach ensures that the task-specific performance is reflective of the data the clients are primarily working with, making the evaluation more aligned with the real-world distribution of data. When analyzing the results for $\alpha=0.1$, FLEx (37.30) ranks closely behind the top method, MoE + FedAvgM (37.41). It is important to note that as a personalized federated learning approach, FLEx is evaluated on each client’s personalized checkpoint, rather than a single global model. This distinction influences its performance relative to methods optimized solely for global averages. The relatively close performance values observed for $\alpha=1.0$ in this experiment are characteristic of the Dirichlet distribution; higher $\alpha$ values in this distribution lead to more balanced data heterogeneity across clients, which tends to reduce the performance gap between different FL methods compared to pathological non-IID scenarios. The key takeaway is that FLEx remains competitive even in scenarios with less extreme data heterogeneity. Its ability to effectively balance personalization with adaptation to a broader data distribution underscores its robustness and adaptability to diverse client needs, standing out from methods focused solely on global performance.
5.4 Performance under LLM-as-a-Judge Evaluation
@@
FLEx (Ours) 6.360 7.993
Figure 2: The figure on the left visualizes the distribution and central tendency of Helpfulness and Harmlessness scores for different methods (Original Model, FLEx, Local). Scatter points and kernel density contours represent the data distribution, with larger markers indicating the average scores. The table on the right presents the average Helpfulness and Harmlessness scores across evaluated methods, demonstrating performance in these two key aspects of open-ended generation. FLEx (Ours) achieves the highest average scores in both Helpfulness (6.360) and Harmlessness (7.993), showing significant improvement, particularly in harmlessness.
In this section, we evaluate the performance of FLEx on an open-ended benchmark Vicuna from helpfulness and harmlessness in two aspects. This experiment was conducted on a set of 20 clients under an IID (Independent and Identically Distributed) scenario to assess the performance of FLEx in open-ended generation tasks, specifically focusing on the aspects of Helpfulness and Harmlessness. The goal of this experiment was to evaluate the model’s ability to maintain high-quality, useful, and safe responses in a more balanced and homogeneous data setting. This IID setting allows for a more straightforward comparison of the different methods under conditions where the data distribution is uniform across clients, ensuring that the performance differences observed are primarily due to the methods themselves rather than data heterogeneity.
The left side of Figure 2 visualizes the distribution of Helpfulness and Harmlessness scores for Original Model, FLEx and Local. The scatter points represent individual evaluations, while the kernel density contours show the overall distribution of scores. The larger markers indicate the average scores for each method. We can see that FLEx not only achieves higher average scores but also exhibits a more concentrated and consistent performance, especially in the Harmlessness dimension. From the table on the right, we observe that FLEx outperforms all other methods in both Helpfulness (6.360) and Harmlessness (7.993), achieving the highest scores in both categories. This is particularly notable in the Harmlessness aspect, where FLEx significantly exceeds all other methods, including MoE + FedAdagrad (the second-best method).
These results demonstrate the advantages of FLEx in open-ended generation tasks, particularly in capturing the subtle distinctions of helpfulness and harmlessness. Compared to other methods, FLEx, through its personalized client models, is able to more finely capture the nuanced features of helpful and harmless content. In terms of helpfulness, FLEx performs exceptionally well, generating content that is not only valuable to users but also accurate and effective. In terms of harmlessness, FLEx stands out especially, as it effectively avoids generating potentially harmful or unsafe outputs, significantly surpassing other methods.
5.5 Communication Cost Analysis
We compare the communication cost of FLEx on MoE-based LLMs. Specifically, we examine the communication cost when applying the LoRA adapter to all MoE experts and transmitting all the experts across clients on Qwen1.5-MoE-A2.7B. The traditional FL cost is approximately 3.41% of the total parameters (i.e., 889M), while the FLEx cost is nearly 0.0659% (i.e., 17.2M) per communication round. It is evident that FLEx uses significantly lower communication costs to achieve high performance.
5.6 Ablation Study
Table 4: Ablation Study Results. ✓ indicates the component is included, and ✗ indicates it is excluded. Here, "No Personalized" denotes that each client prunes the expert using the C4 dataset, a pre-training dataset that does not contain personalized data. Conversely, "Personalized" refers to pruning the expert using the client’s own data.
@@ 5.6 Ablation Study
Pruning Personalized Adaptive Gate Rouge-L
41.17
5.16
7.06
42.67
43.14
In this ablation study, we evaluate the contributions of the three core components of the FLEx framework: Pruning, Personalized Pruning, and Adaptive Gate. Experiments were conducted using the Databricks-dolly-15k dataset with FedAvg as the aggregation method and performance evaluated using the ROUGE-L metric. Results are summarized in Table 4. The baseline model, without any components enabled, achieved a ROUGE-L score of 41.17. Notably, enabling only Pruning drastically reduced the performance to 5.16, underscoring the critical necessity of personalization. Incorporating Personalized Pruning alongside Pruning slightly improved the score to 7.06, yet the performance remained limited, demonstrating that personalization alone is insufficient without effectively leveraging client-specific expertise. Enabling the Adaptive Gate raised performance substantially to 42.67, surpassing the baseline. Finally, integrating all three components—Pruning, Personalized Pruning, and Adaptive Gate yielded the highest performance of 43.14. This result underscores the critical role personalization plays within the FLEx framework, highlighting that its full potential is best realized when combined with the dynamic selection capability of adaptive gating to optimally accommodate diverse client datasets.
@@ 6 Conclusion
In this paper, we introduced FLEx, a federated learning framework tailored for MoE-based large language models. FLEx reduces communication and computation costs by allowing each client to fine-tune and retain a single, local expert while aggregating only the necessary updates from multi-head attention layers. This design effectively balances local personalization with global knowledge sharing, achieving up to 50$\times$ reduction in communication overhead compared to traditional methods. Extensive evaluations show that FLEx outperforms existing federated learning baselines on non-IID tasks, demonstrating its potential for efficient and scalable deployment of MoE-based LLMs in privacy-sensitive settings. Future work will focus on improving scalability and robustness across diverse applications.
@@ Acknowledgement
This work was supported by National Natural Science Foundation of China (62303319, 62406195), Shanghai Local College Capacity Building Program (23010503100), ShanghaiTech AI4S Initiative SHTAI4S202404, HPC Platform of ShanghaiTech University, Core Facility Platform of Computer Science and Communication of ShanghaiTech University, and MoE Key Laboratory of Intelligent Perception and Human-Machine Collaboration (ShanghaiTech University), and Shanghai Engineering Research Center of Intelligent Vision and Imaging.

### v2

@@ 5 Experiments
This section presents comprehensive experiments designed to evaluate the effectiveness of FLEx.
@@ 5.1 Experimental Setup
We utilize the Databricks-dolly-15k [^1] https://huggingface.co/datasets/databricks/databricks-dolly-15k dataset, an open-source collection of instruction-following records. To simulate a non-IID data distribution, we partition the dataset by selecting four subtasks with objective, verifiable answers: classification, closed QA, information extraction, and summarization. We conduct our experiments in a cross-silo federated learning setting to evaluate the generalization performance of FLEx across diverse data distributions. This targeted selection ensures that the data on each client corresponds to a specific task category.
Base Model and Fine-tuning Strategy.
Our experiments use the Qwen1.5-MoE-A2.7B [^2] https://huggingface.co/Qwen/Qwen1.5-MoE-A2.7B model as the backbone. For efficient fine-tuning, we apply LoRA (Hu et al., 2022) to the model’s trainable layers, configured with a rank of 32 and an alpha of 64. The local model on each client is trained using the Adam optimizer with a learning rate of $4\times 10^{-5}$ and a batch size of 4. We use bfloat16 for mixed-precision training and set the maximum input sequence length to 2048. All models are prompted with the Alpaca template. The experiments are conducted on NVIDIA H20 GPUs, each with 96GB of memory.
Federated Training Protocol and Baselines.
We implement the federated learning process using the OpenFedLLM framework (Ye et al., 2024). The training proceeds for 100 communication rounds, with each client processing 5 local steps per round. We compare our proposed method, FLEx, against several established federated learning algorithms: FedAvg (McMahan et al., 2017), FedProx (Li et al., 2020), SCAFFOLD (Karimireddy et al., 2020), FedAvgM (Hsu et al., 2019), FedAdagrad (Reddi et al., 2021), FedYogi (Reddi et al., 2021) and FedAdam (Reddi et al., 2021). For all baseline methods, we apply the respective aggregation algorithm to all trainable parameters of the MoE model, including both the expert and non-expert layers. This represents a standard application of these FL algorithms to the MoE architecture.
Evaluation Metric.
To measure the instruction-following capability of the trained models, we use the ROUGE-L score as the primary evaluation metric. ROUGE-L is well-suited for our chosen tasks as they have objective, verifiable answers (details in Appendix C.9). For fair comparison, we evaluate the checkpoints from the same communication round across all methods.
Table 1: ROUGE-L performance on the Databricks-dolly-15k dataset under a pathological non-IID setting (one task per client). We also report average MMLU scores to assess general knowledge retention. Best and second-best results among FL methods are bolded and underlined, respectively. CLF: Classification, CQA: Closed QA, IE: Information Extraction, Summ: Summarization.
@@
[…]
5.2 Main Results on Instruction Following
Performance under Pathological Non-IID Data Partition.
To test the model’s ability to handle extreme data heterogeneity, we first evaluate its performance in a pathological non-IID setting where each client’s data is confined to a single, distinct task. As shown in Table 1, FLEx achieves the highest average ROUGE-L score, outperforming the next-best federated method, MoE+FedAvgM. Crucially, the inability of most conventional FL algorithms to outperform isolated local training highlights the inherent challenge of forging a single global model from diverging client objectives in highly heterogeneous environments.
To further investigate FLEx’s effectiveness, we evaluated the general knowledge preservation of the personalized models using the MMLU(Wang et al., 2024a) benchmark. FLEx achieves the highest MMLU score among the compared methods, suggesting it effectively preserves the foundational knowledge of the base model. We attribute this to our design of freezing the pretrained, shared experts, which protects them from conflicting and potentially catastrophic updates from specialized clients. This approach enables personalized experts to specialize in local tasks without diminishing the model’s core capabilities.
Table 2: Average ROUGE-L performance on Databricks-dolly-15k under a Dirichlet distribution with $\alpha=0.1$ (higher heterogeneity) and $\alpha=1.0$ (lower heterogeneity). Best and second-best results are bolded and underlined.
@@ 5.2 Main Results on Instruction Following
[…]
Performance under Dirichlet-based Non-IID Setting.
Finally, we simulate more realistic, yet still challenging, non-IID scenarios by partitioning the Databricks-dolly-15k dataset among 10 clients using a Dirichlet distribution. Table 2 shows that FLEx remains highly competitive. It outperforms the other methods for $\alpha=1.0$, a setting with moderate heterogeneity, and the second-best for $\alpha=0.1$, a setting with higher heterogeneity. While the performance gap between methods narrows as data distributions become more balanced, FLEx’s consistent leading performance highlights its adaptability. This robust performance can be attributed to its effective personalization, which enables the model to generalize across various degrees of data heterogeneity, not just in extreme cases.
Helpfulness and Harmlessness Evaluation.
We assessed open-ended generation quality using the Vicuna benchmark in an IID setting with 20 clients. As shown in Figure 3, FLEx achieves the highest scores for both Helpfulness (6.360) and Harmlessness (7.993). The significant improvement in harmlessness is particularly noteworthy, suggesting that FLEx’s personalized approach is better at capturing the nuances required to generate high-quality, safe, and useful responses, outperforming both baselines and the original pre-trained model.
@@
[…]
Figure 3: Evaluation of Helpfulness and Harmlessness on the Vicuna benchmark. Left: Score distributions for the base model, local training, and FLEx. Larger markers indicate average scores. Right: Average scores for all evaluated methods. FLEx achieves the highest scores on both metrics, with a notable improvement in harmlessness.
5.3 Ablation and Component Analysis
Ablation Study on Core Components.
We conducted an ablation study on the Databricks-dolly-15k dataset under a pathological non-IID setting to evaluate the contributions of our two core components: personalized expert grafting and the adaptive gate. As shown in Table 3, naively grafting an expert without our adaptive gate leads to a catastrophic performance collapse, as shown in settings (b) and (c). This occurs because naively inserting an expert without the adaptive gate destabilizes the model’s internal dynamics. The gate is crucial for learning to appropriately scale and integrate the expert’s contributions. Without it, the raw outputs from the grafted expert can overwhelm or corrupt the knowledge pathways of the frozen base model.
Introducing the adaptive gate, even with a generic expert (d), recovers performance and surpasses the FedAvg baseline. This result highlights the gate’s critical role in dynamically merging the knowledge from the frozen and grafted experts. Finally, our complete FLEx model (e), which combines the adaptive gate with a personalized expert, achieves the best results. The improvement from (d) to (e) demonstrates the distinct benefit of tailoring the grafted expert to a client’s local data. These results confirm that both components are important: the adaptive gate enables effective knowledge integration, and personalization enhances local adaptation.
Table 3: Ablation study on the core components of FLEx. We analyze the effects of the expert grafting criterion (personalized vs. generic) and the adaptive gating mechanism.
Model Configuration ROUGE-L
(a) Baseline (FedAvg, no FLEx components) 41.17
— Naively Integrating a Grafted Expert —
(b) Generic Grafting w/o Adaptive Gate 5.16
(c) Personalized Grafting w/o Adaptive Gate 7.06
— Integrating via Adaptive Gating —
(d) Full FLEx w/ Generic Grafting 42.67
[…]
Impact of Training and Communication Strategy.
To validate our architectural choices, we conduct an ablation study evaluating the impact of the training strategy (which experts to train) and the communication strategy (which parameters to share). The results in Table 4 lead to two key observations. First, since each client’s experts specialize in distinct local data distributions, naively aggregating all client experts induces destructive interference, which degrades overall performance (row 1 vs. 3 and 2 vs. 4). In contrast, sharing only the common non-expert parameters is more effective, as they capture shared knowledge without this specialized interference. Second, under identical communication costs (rows 3-5), FLEx’s personalized expert grafting (row 5) substantially outperforms training all experts (row 3) or only the most activated one (row 4). These findings confirm that FLEx’s strength lies in its effective combination of targeted personalization and communication efficiency, rather than simply concentrating training resources.
Table 4: Performance comparison of different strategies for training and sharing MoE modules on Databricks-dolly-15k (pathological non-IID). Our method (FLEx) achieves the highest performance with the lowest communication cost. CP: Communication Parameters, TP: Trainable Parameters.
Strategy Databricks-dolly-15k Cost (%)
Trained Experts Shared Parameters CLF CQA IE Summ Avg CP $\downarrow$ TP $\downarrow$
All Non-Experts + All Experts 51.39 34.52 37.23 41.54 41.17 3.4172 3.4172
Activated Non-Experts + All Experts 50.83 34.34 36.26 41.59 40.75 0.1327 0.1327
All Non-Experts Only 51.10 32.51 41.39 40.60 41.40 0.0672 3.4172
Activated Non-Experts Only 51.48 34.77 37.23 41.74 41.30 0.0672 0.1327
[…]
Ablation on the Gating Activation Function.
We investigated the choice of activation function for the grafted expert’s gating mechanism, as defined in Eq. (9). We compared our default sigmoid function against two common alternatives: ReLU and Tanh. The results, summarized in Table 5, demonstrate that the sigmoid-based gate consistently and significantly outperforms the other functions. A detailed discussion on the rationale behind this performance difference is shown in Appendix C.3.
@@ 5.3 Ablation and Component Analysis
Method CLF CQA IE Summ Avg
ReLU 51.16 33.02 37.51 41.78 40.86
Tanh 41.40 20.40 31.34 37.88 32.75
Sigmoid 52.10 38.23 40.13 42.09 43.13
Compatibility with PEFT.
FLEx is compatible with Parameter-Efficient Fine-Tuning (PEFT) methods like LoRA. By applying LoRA or other PEFT methods to the trainable components in FLEx, only a small set of adapter parameters are updated instead of the full weights. This strategy leverages dynamic sparsity to further reduce both computational and communication overhead in FL.
@@ 6 Conclusion
In this work, we introduced FLEx, a novel federated learning framework designed to effectively harness the power of pretrained MoE-based LLMs in settings with heterogeneous data. Our approach addresses the critical challenges of catastrophic forgetting and prohibitive communication costs by selectively aggregating only the shared, dense parameters while freezing the pretrained experts to preserve their specialized knowledge. For personalization, we proposed a novel expert grafting mechanism that allows clients to construct lightweight, customized experts tailored to their local data, which are then dynamically integrated with the frozen experts via a jointly trained gating mechanism. Extensive evaluations demonstrate that FLEx significantly outperforms established federated baselines on personalization tasks across diverse non-IID datasets, while simultaneously preserving the model’s general knowledge as validated by strong performance on the MMLU benchmark.
@@ Acknowledgement
This work was supported by National Natural Science Foundation of China (62303319, 62406195), Shanghai Local College Capacity Building Program (23010503100), ShanghaiTech AI4S Initiative SHTAI4S202404, HPC Platform of ShanghaiTech University, and MoE Key Laboratory of Intelligent Perception and Human-Machine Collaboration (ShanghaiTech University), Shanghai Engineering Research Center of Intelligent Vision and Imaging. This work was also supported in part by computational resources provided by Fcloud CO., LTD.
Reproducibility statement
To support the reproducibility of our research, we detail our experimental setup and hyperparameters in Section 5.1. All datasets used are publicly available. Furthermore, our source code will be publicly released upon publication.
Ethics statement
This research was conducted ethically. We used only publicly available datasets, ensuring that no sensitive or private data was involved. We have also considered the broader impact of this work and foresee no negative societal consequences.